OpenStudy (anonymous):

a data set consists of the following data points (3,5),(5,8),(6,13) the slope of best fit line is 2.5. find the y intercept of this line

OpenStudy (anonymous):

really need help

OpenStudy (anonymous):

i think i -2.5???

OpenStudy (anonymous):

y = m x + b y = 2.5 x + b 5 = 2.5 * 3 + b 5 = 7.5 + b b = - 2.5 y = 2.5 x + b 8 = 2.5 * 5 + b 8 = 12.5 + b b = - 4.5 y = 2.5 x + b 13 = 2.5 * 6 + b 13 = 15 + b b = -2 Averaging the b's, ( - 2.5 - 4.5 - 2) / 3 = -9/3 = -3.

OpenStudy (anonymous):

so its -3

OpenStudy (anonymous):

it was thank you so much seeing it worked out help me out alot :)

OpenStudy (anonymous):

You're welcome. :)